OpenStudy (anonymous):

simplify -(-8m)-(-7m+2)

OpenStudy (zzr0ck3r):

8m +7m -2 15m-2

OpenStudy (zzr0ck3r):

-*(-) = +

OpenStudy (anonymous):

i was getting 15m+2

OpenStudy (zzr0ck3r):

-(7m+2) = -7m-2

OpenStudy (zzr0ck3r):

the minus sign gets distributed to all terms in the ()

OpenStudy (anonymous):

oh ok ;) thanks

OpenStudy (zzr0ck3r):

make sence?

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

so -(-a+b-(-c)-d) = a-b-c+d